OpenStudy (anonymous):

use substitution to solve the system {12x +y=21 , 18x-y=-36

OpenStudy (anonymous):

Make the substitution y= 18x+36 from equation 2. Sub it into equation 1. 12x + (18x+36) = 21 30x = 21 - 36 >> x = -15/30= -1/2 Now solve for y, y = 18(-1/2) + 36 = 27 x = -1/2, y = 27